OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 4 sin2 x - 4 sin x + 1 = 0

OpenStudy (campbell_st):

its basically a quadratic equation that can be factorised (2sin(x) -1)^2 = 0 so you need to solve 2sin(x) -1 = 0 consider the domain when looking for solutions.. as there are 2 answers.

OpenStudy (anonymous):

so sin x = 1/2? I'm still confused as to how you got there...

OpenStudy (anonymous):

@campbell_st

OpenStudy (loser66):

let t= sinx , so you have sin^2x = t^2, right? replace to the equation, you have 4t^2 -4t +1=0, solve for t, then plug back to t = sinx , solve for x , don't forget interval

OpenStudy (anonymous):

Oh! That makes sense. Thank you!